A particle is moving with a uniform acceleration along a straight line. Its velocities at A and B are respectively 7m/s and 17m/s^. M is the mid point of A and B. If t1is the time tajen to go from A to M and t2 is the time taken to go from M to B. Then what would be the ratio of t1/t2? 
It’s answer is 3:2 but I am not getting the solution May you please help me out...

Grade:11

2 Answers

Arun
3 years ago
given conditions in the question are
particle is having uniform velocity.
velocity at A is 7m/s
velocity at B is 17m/s
M is midpoint of AB. so let us consider the distance between AB is “d”
therefore distance between AM is d/2 and distance between MB is d/2
here t1 is the time taken to reach A to M. 
so velocity=displscement /time
time= displacement / velocity
for case AM
t1 =d/2*7 (*means multiplication)
similarily t2=d/2*17
t1/t2= 17/7 (on solving)
therefore the solution is t1/t2= 17/7